# Towards Explainable Bilingual Multimodal Misinformation Detection and Localization

### v1

@@ 6 Conclusion and Limitation
Conclusion. We study the task of detecting multimodal inconsistencies in news images with bilingual (Chinese-English) subtitles. To support this, we introduce BiMiBench, a large-scale benchmark with 104K samples featuring realistic visual edits and cross-lingual mismatches. We propose BiMi, a bilingual multimodal framework combining a three-stage post-training strategy and retrieval-augmented reasoning. Experiments show that BiMi achieves state-of-the-art performance in classification, localization, and explanation, highlighting its effectiveness in real-world scenarios.
Limitation and future work. While BiMi achieves strong results on bilingual multimodal detection, several limitations remain. First, the current setup focuses on Chinese-English subtitles; extending to more languages would enhance generalization. Second, localization is limited to bounding boxes, which may miss fine-grained manipulations—future work could explore pixel-level segmentation. Third, BiMiBench samples use fixed subtitle styles and editing templates; expanding visual diversity and subtitle formats would better reflect real-world cases. Finally, while BiMi offers solid performance, its reasoning and explanation ability can be further improved, especially in ambiguous scenarios.
Appendix
Section A provides details on the dataset, including: dataset statistics, misinformation generation strategy, example visualization, and manual verification.
Section B outlines implementation and experimental details, including: model architecture and training settings, hyperparameters.
Section C discusses the broader impact and ethical implications of our work.
Section D presents qualitative examples.
Section E analyzes the failure case.
Appendix A Dataset Details
Dataset Statistics. BiMiBench comprises 104,000 samples spanning six fine-grained misinformation categories. Figure 7 summarizes key dataset properties. The left panel shows the distribution of samples across categories, ensuring diverse coverage of manipulation types.
@@ Appendix A Dataset Details
Misinformation Generation Strategy. To simulate realistic multimodal misinformation, we generate manipulated subtitles by introducing subtle yet misleading semantic alterations. These edits preserve grammatical fluency and contextual plausibility while modifying key factual elements such as subjects, causal relations, locations, or numerical values. We further introduce bilingual inconsistencies by translating the manipulated English subtitles into Chinese, allowing semantic shifts to emerge naturally during translation. This strategy enables diverse and challenging misinformation scenarios that reflect real-world cross-modal and cross-lingual manipulations. We design task-specific prompts to guide misinformation generation, ensuring fluency and semantic plausibility while introducing factual inconsistencies; an example prompt used in this process is shown in Table 8.
<system> Based on the image, create a fictional, sensational news headline or summary in one sentence. The story must be clearly fake but sound plausible, and must be no more than 30 words. The news should sound outrageous, scandalous, or unbelievable, yet still written in a serious, news-like tone. </system>
@@
Example Visualization. Figure 8 illustrates representative BiMiBench samples, covering various manipulation types, language inconsistencies, and corresponding ground truth labels.
@@
Manual Verification. To ensure data quality and annotation reliability, all samples in BiMiBench were manually verified by a team of five trained annotators, each with backgrounds in journalism, linguistics, or media studies, and prior experience with fact-checking or misinformation analysis. Annotators received targeted instruction on identifying realistic image manipulations, assessing semantic consistency across English-Chinese subtitle pairs, and applying fine-grained manipulation categories. During the review process, they checked the accuracy of tampered regions, validated bounding box alignment, and evaluated bilingual subtitles for cross-lingual fidelity and plausibility. Samples with OCR errors, ambiguous edits, mistranslations, or low visual quality were discarded. This multi-layered, expert-driven verification process ensures that BiMiBench maintains high annotation quality, semantic precision, and supports reliable benchmarking for multilingual multimodal misinformation detection.
Appendix B Implementation Details
@@
Model Architecture. Our framework is built upon the publicly released Gemma 3 multimodal language model. The vision encoder is initialized from a pretrained ViT-L model, and remains frozen throughout all training stages. LoRA parameters are fine-tuned. The model accepts the inputs of the image with English and Chinese subtitle, which serialized into a unified prompt format. The image is encoded through a unified visual-language backbone.
@@
Tokenization and Input Format. Both English and Chinese subtitles are tokenized using the SentencePiece tokenizer aligned with the Gemma 3 backbone. The input is formatted into structured prompts that specify the prediction task, marking different modalities. Input length is capped at 512 tokens.
Hyperparameter Summary. We fine-tune BiMi using the Gemma 3 model (4B) with a batch size of 8, a learning rate of $1\times 10^{-5}$, and a maximum input length of 512 tokens. The model is optimized with AdamW, using $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a weight decay of 0.01. To stabilize training, we use gradient accumulation (4 steps) and mixed precision (FP16). The GRPO reward coefficient $\lambda$s are all set to 1. These hyperparameters are kept consistent across all three training stages unless otherwise specified.
@@
<system> You are an expert in misinformation detection area. A conversation between User and Assistant. The user asks a question, and the Assistant solves it. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think><answer> answer here </answer>. Assign one of the following six categories based on their mutual alignment: All Consistent: The image aligns well with both Chinese and English captions. No signs of manipulation or misalignment. Image Manipulated: The image is manipulated. Both captions truthfully describe the manipulated image. Both Misaligned: Both Chinese and English captions are manipulated. Neither caption correctly describes the image. Chinese Misaligned: Only the Chinese caption is manipulated. The English caption aligns correctly with the image. English Misaligned: Only the English caption is manipulated. The Chinese caption aligns correctly with the image. All Inconsistent: The image, Chinese caption, and English caption are all manipulated and mutually inconsistent. </system> <user> Please analyze the given image containing both Chinese and English subtitles and complete the following three tasks: (1) Classification Task: classify the alignment between the image and the subtitles into one of the following six categories: "all consistent", "image manipulated", "both subtitles misaligned with image", "only English aligned", "only Chinese aligned", "all inconsistent". (2) Manipulation Detection: if the image has been manipulated, return one or more bounding boxes for the manipulated regions in the format: {"bbox":[x_min, y_min, x_max, y_max]}. If no manipulation is found, return an empty list. (3) Decision Explanation: briefly explain your thinking before the classification and any detected regions. Return your output using the following format, wrapped in tags: <think> Your explanation here. </think> <answer> "classification": "result", "region": [{"bbox": [x_min, y_min, x_max, y_max]}] </answer> </user>
@@
Appendix C Broader Impact and Ethics
BiMi is designed to enhance the interpretability and reliability of multimodal misinformation detection, particularly in the context of bilingual news media. By providing both classification outputs and natural language explanations, BiMi promotes transparency in AI decision-making and supports fact-checkers, journalists, and platform moderators in identifying subtle manipulations across images and subtitles. To ensure ethical standards, all data samples in BiMiBench have been manually reviewed, linguistic clarity, and annotation consistency. Low-quality, ambiguous content has been excluded. The dataset construction avoids personally identifiable information and sensitive content.
BiMi is detection-focused and does not produce or promote fake content, reducing the risk of misuse. Nonetheless, we recognize that any publicly released model or dataset could potentially be repurposed. To mitigate this, we include usage guidelines and encourage responsible deployment, especially in downstream applications that may involve high-stakes decisions. We believe this work contributes positively by advancing research on robust, interpretable AI tools for combating harmful information while maintaining strong ethical safeguards throughout the development and release process.
Appendix D Qualitative Examples
Figure 9 10 11 12 13 14 15 16 show the examples.
Figure 9: Case of All Consistency.
@@ Appendix D Qualitative Examples
Figure 11: Case of Image Manipulation.
@@
Appendix E Failure Cases
Figure 17 shows a failure case where the model misclassified an image manipulation example as “All Inconsistent.” The English caption aligns with the image, but the Chinese one introduces a false espionage narrative. Key issues include: misclassification; incorrect caption judgment—only the Chinese caption was wrong; missing bounding box. This highlights weaknesses in manipulation detection, multilingual judgment, and grounding.
Figure 17: Failure case

### v3

@@ 6 Conclusion and Limitation
Conclusion. We study the task of detecting multimodal inconsistencies in news images with bilingual subtitles. To support this, we introduce BiMiBench, a large-scale benchmark with 104K samples featuring realistic visual edits and cross-lingual mismatches. We propose BiMi, a bilingual multimodal framework combining a three-stage post-training strategy and retrieval-augmented reasoning. Experiments show that BiMi achieves state-of-the-art performance across tasks, highlighting its effectiveness in real-world scenarios. This work advances research on bilingual and multimodal misinformation detection.
Limitation. While BiMi achieves strong results on bilingual multimodal detection, several limitations remain. First, it currently targets only Chinese–English subtitles; extending to additional languages would improve cross-lingual generalization. Second, localization is restricted to bounding boxes and may miss fine-grained manipulations. Third, the data diversity of BiMiBench is limited: beyond broader subtitle templates and editing styles, richer sources spanning multiple languages and topical domains are needed to better reflect real-world variability. Fourth, retrieval introduces extra latency and depends on external sources, which can occasionally return noisy or outdated information. Finally, BiMi’s reasoning capability could be strengthened, especially for cases with conflicting multimodal evidence.
Overview of Supplementary Material
Section A: Dataset and ethics.
Section B: Additional experiments and cases.
Section C: Implementation details.
Appendix A Dataset Construction and Quality Control
A.1 Dataset Statistics
BiMiBench comprises 104,000 samples spanning six fine-grained misinformation categories. Figure 6 summarizes the key dataset properties. Beyond the balanced distribution across categories (left), BiMiBench exhibits substantial diversity in both visual content and bilingual subtitle structures. The images cover a wide range of news domains, including politics, disasters, sports, daily life, technology, and international affairs, ensuring that manipulation patterns are not confined to a narrow set of scenes. The subtitles vary in length, writing style, and translation quality across topics in both Chinese and English. Such multimodal richness helps simulate real-world bilingual misinformation scenarios and provides a challenging, comprehensive testbed for fine-grained manipulation detection.
@@ A.1 Dataset Statistics
A.2 Misinformation Generation Strategy.
To simulate realistic multimodal misinformation, we generate manipulated subtitles by introducing subtle yet misleading semantic alterations. These edits preserve grammatical fluency and contextual plausibility while modifying key factual elements such as subjects, causal relations, locations, or numerical values. We further introduce bilingual inconsistencies by translating the manipulated English subtitles into Chinese, allowing semantic shifts to emerge naturally during translation. This strategy enables diverse and challenging misinformation scenarios that reflect real-world cross-modal and cross-lingual manipulations. We design task-specific prompts to guide misinformation generation, ensuring fluency and semantic plausibility while introducing factual inconsistencies; an example prompt used in this process is shown in Table 7.
@@
<<system>> Based on the image, create a fictional, sensational news headline or summary in one sentence. The story must be clearly fake but sound plausible, and must be no more than 30 words. The news should sound outrageous, scandalous, or unbelievable, yet still written in a serious, news-like tone. <</system>>
Data Diversity. To ensure broad and representative diversity in the synthetic misinformation, we introduce variation at three complementary stages. Data source: we begin with the large and professionally curated VisualNews corpus, whose wide topical coverage—politics, economics, sports, culture—naturally provides a variety of writing styles and visual contexts, ensuring that manipulations are not limited to a narrow set of news domains. Editing: on top of this diverse base we generate multiple kinds of factual distortions, ranging from single–word entity swaps to complex changes in temporal references, numerical values and causal relations. We also randomize prompt formulations and the sampling parameters of the generation model so that the manipulated bilingual subtitles exhibit heterogeneous and unpredictable inconsistencies. Label assignment: after generation, each candidate sample is manually reviewed by trained annotators and assigned to one of the six defined misinformation categories, reflecting the precise combination of visual edits and cross-lingual subtitle inconsistencies. This three-stage process guarantees that BiMiBench covers a wide spectrum of realistic manipulation patterns and prevents overfitting to any type of false information.
A.3 Review Protocol
To ensure data quality and annotation reliability, all samples in BiMiBench were manually verified by a team of five trained annotators, each with backgrounds in journalism, linguistics, or media studies, and prior experience with fact-checking or misinformation analysis. Annotators received targeted instruction on identifying realistic image manipulations, assessing semantic consistency across English-Chinese subtitle pairs, and applying fine-grained manipulation categories. During the review process, they checked the accuracy of tampered regions, validated bounding box alignment, and evaluated bilingual subtitles for cross-lingual fidelity and plausibility. Samples with OCR errors, ambiguous edits, mistranslations, or low visual quality were discarded. This multi-layered, expert-driven verification process ensures that BiMiBench maintains high annotation quality, semantic precision, and supports reliable benchmarking for multilingual multimodal misinformation detection.
A.4 Ethical and Licensing Considerations
BiMiBench is derived from the publicly available VisualNews corpus, which provides professionally curated image–text news pairs under a license permitting academic research. All original images and captions remain the intellectual property of their respective news organizations; our release is strictly for non-commercial, research purposes and requires users to acknowledge the original sources and this benchmark in any derivative work.
While BiMiBench is designed to advance scientific understanding of multimodal misinformation detection, it necessarily contains intentionally manipulated content: images edited to introduce subtle visual changes and bilingual subtitles rewritten to create semantic inconsistencies. This “dual nature”—a resource built for research, yet embedding realistic examples of fabricated content—poses an inherent risk of misuse if taken out of context or circulated without clear academic framing.
To mitigate these risks, we (i) clearly mark all manipulated samples as synthetic, (ii) provide the data only for research and educational use, and (iii) require that any redistribution or downstream work include appropriate citations and comply with relevant copyright and data-protection regulations. We explicitly prohibit using BiMiBench to create or disseminate deceptive media outside controlled research settings. These safeguards aim to ensure that the benchmark remains a tool for studying and countering misinformation rather than a source of it.
A.5 Example Visualization.
Figure 7 illustrates representative BiMiBench samples, showcasing a wide range of manipulation patterns, cross-lingual inconsistencies, and subtitle complexities.
@@ A.4 Ethical and Licensing Considerations
Appendix B Additional Experiments
B.1 Reference Generation
To construct high-quality reference explanations for automatic evaluation, we adopt a two-stage human–LLM pipeline. First, for each sample we prompt GPT-4o with task-specific, label-grounded templates to generate two candidate explanations conditioned on the manipulated modality and misinformation type. Second, a team of five trained annotators—each with a background in journalism, media studies, or NLP—independently reviewed the two candidates and selected the one that best satisfied five ground-truth–aligned dimensions: (i) Subtitle alignment, (ii) Visual detail consistency, (iii) Reasoning consistency, (iv) Clarity and readability, and (v) Completeness. The chosen explanation for each sample was retained as the pseudo-ground truth used in BERTScore evaluation. This procedure ensures that every reference explanation is both fluent and factually faithful while keeping the generation process efficient and reproducible.
Table 8: Human evaluation criteria and scoring scheme for selecting reference explanations. Annotators scored each criterion on a 0–2 scale (0 = not satisfied, 1 = partially satisfied, 2 = fully satisfied), and explanations with a total score below 7 were discarded or revised.
Criterion Description Score
Subtitle alignment Checks consistency between English–Chinese subtitles and the described manipulation. 0–2
Visual detail consistency Ensures the explanation correctly reflects visible tampered regions or edits. 0–2
Reasoning consistency Evaluates logical coherence between detected evidence and the explanation. 0–2
Clarity and readability Measures fluency and conciseness suitable for news-style explanations. 0–2
Completeness Ensures all manipulated elements are covered without omitting key evidence. 0–2
B.2 Performance on Real-World Samples
To assess real-world generalization, we manually collected 100 recent social-media posts spanning political, health, and international news. These posts contain image–caption manipulations, cross-lingual inconsistencies, and subtle visual edits typical of in-the-wild misinformation. Because the authenticity of such posts cannot be verified with ground-truth labels, the 100 examples were expert-curated: five annotators with backgrounds in journalism and fact-checking jointly reviewed the content and reached consensus on whether each case constitutes misinformation. This human consensus served as the reference label set.
Evaluated against these expert judgments, BiMi correctly identified 81 of the 100 curated cases (81% accuracy) . These findings suggest that BiMi can transfer effectively from synthetic benchmarks to uncontrolled real-world conditions, even when the ground truth is based on expert assessment rather than definitive factual verification.
We analyze a real-world social media post with subtle bilingual inconsistency (Figure 8). The removed ”student” in the Chinese translation alters the meaning. Without retrieval, the model misclassifies due to a lack of context. The retrieval result provides a key clue for classification. BiMi correctly identifies the inconsistency, showing its benefit for nuanced cross-lingual reasoning.
Figure 8: Social media example with subtle bilingual inconsistency. The Chinese translation omits ”student”, changing the meaning and potentially misleading Chinese readers. Middle: InternVL3 result; Right: BiMi result.
B.3 Effect of OCR Module
This section provides brief supplementary observations expanding on the OCR analysis in the main paper. While the core quantitative results are already reported, here we include additional qualitative notes.
During manual inspection, we found that OCR errors most commonly occur in Chinese subtitles containing long sentence structures or low-resolution text, which aligns with the degradation patterns shown in the main paper. In contrast, English subtitles are generally recognized cleanly, leading to smaller performance differences between ground-truth and OCR-extracted inputs. In many real-world samples, OCR inaccuracies do not alter the final classification category but may affect explanation specificity. We therefore recommend simple preprocessing steps—such as cropping subtitle regions or modest contrast enhancement—to stabilize OCR quality when applying the model outside the benchmark setting.
These findings show that BiMi remains largely robust to moderate OCR noise; the observed degradation originates mainly from preprocessing rather than from limitations in the model’s bilingual reasoning. When clean Chinese subtitles are provided, BiMi’s accuracy approaches the English benchmark. Improving OCR fidelity can enhance overall system reliability in real-world deployments.
B.4 Analysis of Bilingual Performance Gap
BiMi achieves higher accuracy on English-only inputs (82.48% ACC) than on Chinese-only inputs (72.32% ACC). To disentangle the causes of this gap, we conducted a controlled evaluation using ground-truth subtitles versus OCR-extracted subtitles, as summarized in Table 9.
Table 9: Impact of OCR on classification accuracy.
Language GT (%) OCR (%) Drop (%)
English 84.2 83.0 -1.2
Chinese 80.0 75.3 -4.7
The results show that Chinese performance is far more sensitive to OCR noise, with a 4.7% drop compared to only 1.2% for English. Thus, OCR quality in Chinese subtitles is the primary factor behind the overall 10% gap.
Secondary contributors include (i) training data imbalance: VisualNews contains more English captions, which biases the backbone model toward English, and (ii) language-model priors: although Gemma 3 supports bilingual reasoning, its representations are stronger for English. When clean Chinese subtitles are provided, BiMi’s accuracy approaches the English benchmark, confirming that OCR errors are the dominant bottleneck.
To close this gap, future work will focus on (1) enhancing Chinese OCR through confidence-based filtering and multi-OCR consensus, (2) exploring LLM-based subtitle repair to recover noisy outputs, and (3) fine-tuning on more balanced bilingual data. Overall, the performance difference is mainly driven by OCR noise, with data imbalance and model priors as secondary factors.
B.5 Computation and Efficiency
BiMi employs a lightweight multimodal backbone (Gemma 3–4B) to maintain manageable training and inference costs. All experiments are conducted on a single NVIDIA V100 48 GB GPU. End-to-end inference latency is determined by three components: the model forward pass, OCR-based subtitle extraction, and the external retrieval module. Without retrieval, pure model inference averages 320 ms per sample. Enabling the full retrieval pipeline—including OCR and Google Search API queries—adds approximately 0.8–1.2 s of overhead, depending on network conditions and hardware. This additional cost remains acceptable for offline evaluation and is broadly consistent with other retrieval-augmented MLLMs.
To avoid retrieval becoming a single point of failure, BiMi incorporates a graceful-degradation mechanism: if OCR or web search times out or returns no relevant context, the model skips the retrieved evidence and falls back to pure visual–textual reasoning, keeping classification and explanation performance stable. During training, the vision encoder is frozen and only LoRA adapters are updated, which keeps GPU memory usage below 40 GB and allows a batch size of 16. These design choices ensure that BiMi can be reproduced on standard academic hardware while maintaining reasonable inference latency and memory footprint.
B.6 Failure Case Study
Figure 9 shows a failure case where the model misclassified an image manipulation example as “All Inconsistent.” The English caption aligns with the image, but the Chinese one introduces a false espionage narrative. Key issues include (1) misclassification, (2) incorrect caption judgment (only the Chinese caption was wrong), and (3) a missing bounding box.
Figure 9: Failure case
B.7 Qualitative Cases
Figures 10, 11, 12, 13, 14, 15, 16, 17 show the examples.
Appendix C Implementation Details
C.1 Training
@@
Model Architecture. Our framework is built upon the publicly released Gemma 3 multimodal language model. The vision encoder is initialized from a pretrained ViT-L and kept frozen in all training stages, while LoRA adapters are fine-tuned on top of the language backbone to reduce computational cost. The model accepts as input the news image with its English–Chinese subtitles, serialized into a unified prompt that embeds the subtitles as text tokens and the image as patch-level embeddings. Visual tokens from the frozen encoder are projected into the language model’s embedding space and fused with textual tokens through the cross-modal attention layers of Gemma 3, enabling joint reasoning over visual regions and bilingual text. This design allows efficient adaptation to our task while preserving the strong multilingual and vision–language alignment of the pretrained model.
@@
Tokenization and Input Format. Both English and Chinese subtitles are tokenized using the SentencePiece tokenizer aligned with the Gemma 3 backbone to ensure consistent multilingual encoding. The input is formatted into a structured prompt that specifies the prediction task and clearly marks different modalities; visual features from the frozen ViT-L encoder are projected into the same embedding space and appended as a sequence of visual tokens. The combined text and image tokens are then fed to the cross-modal transformer, with the total input length (text + visual tokens) capped at 512 text tokens plus the visual tokens.
[…]
@@
C.2 Retrieve Module
To improve adaptability to rapidly emerging misinformation, we design a retrieval module that operates only at inference time and does not require additional training. First, we apply OCR to extract both the Chinese and English subtitles embedded in the news image. The two subtitles are concatenated to form a bilingual query, which is then sent to the Google Search API. The API returns up to the top-3 relevant passages; if no relevant documents are found, the retrieval component simply returns an empty string. The retrieved text snippets are concatenated and prepended to the model’s input prompt, yielding a unified representation $\mathcal{P}=\operatorname{concat}(R,I,S)$, where $R$, II and $S$ denote the retrieved context, the image and the bilingual subtitles, respectively. This augmented prompt is fed to the MLLM backbone so that cross-attention layers can jointly encode the external evidence and the visual–textual cues. Because the module only provides auxiliary context and does not participate in training, the final predictions remain primarily grounded in the image–subtitle content while benefiting from timely external information when available.
@@ C.2 Retrieve Module
<<system>> You are an expert in misinformation detection area. A conversation between User and Assistant. The user asks a question, and the Assistant solves it. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <<think>> <</think>> and <<answer>> <</answer>> tags, respectively, i.e., <<think>> reasoning process here <</think>><<answer>> answer here <</answer>>. Assign one of the following six categories based on their mutual alignment: All Consistent: The image aligns well with both Chinese and English captions. No signs of manipulation or misalignment. Image Manipulated: The image is manipulated. Both captions truthfully describe the manipulated image. Both Misaligned: Both Chinese and English captions are manipulated. Neither caption correctly describes the image. Chinese Misaligned: Only the Chinese caption is manipulated. The English caption aligns correctly with the image. English Misaligned: Only the English caption is manipulated. The Chinese caption aligns correctly with the image. All Inconsistent: The image, Chinese caption, and English caption are all manipulated and mutually inconsistent. <</system>> <<user>> Please analyze the given image containing both Chinese and English subtitles and complete the following three tasks: (1) Classification Task: classify the alignment between the image and the subtitles into one of the following six categories: "all consistent", "image manipulated", "both subtitles misaligned with image", "only English aligned", "only Chinese aligned", "all inconsistent". (2) Manipulation Detection: if the image has been manipulated, return one or more bounding boxes for the manipulated regions in the format: {"bbox":[x_min, y_min, x_max, y_max]}. If no manipulation is found, return an empty list. (3) Decision Explanation: briefly explain your thinking before the classification and any detected regions. Return your output using the following format, wrapped in tags: <<think>> Your explanation here. <</think>> <<answer>> "classification": "result", "region": [{"bbox": [x_min, y_min, x_max, y_max]}] <</answer>> <</user>>
Figure 10: Case of All Consistent.
@@
Figure 12: Case of Image Manipulated.